OpenStudy (anonymous):

The length of a rectangle is 3 cm more than twice its width. If the perimeter of the rectangle is 54 cm, find the length of the rectangle

OpenStudy (anonymous):

Set up a system of equations. w = width 3+2w = length if the perimeter is 54, use that as the total of all the sides. w+w+3+2w+3+2w = 54 6w+6=54 6w=48 w=8 3+2(8) = 19 so the width is 8 cm and the length is 19 cm.